OpenStudy (anonymous):

Chapter 6:Coordinate Geometry Given that A(6,8),B(p,-3) and C(-4,-6) are three points lying on a straight line such that AB:BC=m:n.Find a)the ratio m:n b)the value of p

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align}& \normalsize \text{For finding 'p' u need to equate the slope .}\hspace{.33em}\\~\\ &\normalsize \text{as the three points are collinear AB and BC should have same slope.}\hspace{.33em}\\~\\ &\dfrac{8-(-3)}{6-p}=\dfrac{-3-(-6)}{p-(-4)} \end{align}}\)

OpenStudy (anonymous):

okay

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{u can find m:n only after finding 'p'}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

\[p=-\frac{ 13 }{ 7 }\]

OpenStudy (mathmath333):

correct

OpenStudy (mathmath333):

For finding \(m:n\) u need the section formula

OpenStudy (mathmath333):



OpenStudy (anonymous):

okay

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} x=\dfrac{nx_1+mx_2}{m+n}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

put the values and get the answer

OpenStudy (anonymous):

\[-\frac{ 13 }{ 7 }=\frac{ 6n-4m }{ m+n }\]

OpenStudy (mathmath333):

\(A(6,8),B(p,-3) ~and~ C(-4,-6)\\ A(x_1,y_1),B(x,y)~and~C(x_2,y_2)\)

OpenStudy (anonymous):

\[-\frac{ 13 }{ 7 }=\frac{ 6n-4m }{ m+n }\]

OpenStudy (mathmath333):

yea simplify it

OpenStudy (anonymous):

\[-13m-13n=42n-28m\]

OpenStudy (anonymous):

\[15m-55n=0\]@mathmath333

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

Am i correct? @mathmath333

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

what should we do next? @mathmath333

OpenStudy (mathmath333):

find \(\dfrac{m}{n}\)

OpenStudy (anonymous):

i'm not sure how to do,can u show me pls?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} 15m-55n=0\hspace{.33em}\\~\\ 3m-11n=0\hspace{.33em}\\~\\ 3m=11n\hspace{.33em}\\~\\ \dfrac{m}{n}=\dfrac{11}{3} \end{align}}\)

OpenStudy (anonymous):

Thank you @mathmath333 :)